OpenStudy (anonymous):

Intuitively, \[\int\limits \limits_{-\infty}^{\infty} x^{3}dx\] should be 0 because it is an odd function, but apparently it diverges. Can anyone give me an intuitive explanation why?

OpenStudy (anonymous):

HOW IT IS ZERO?? I AM GETTING INFINITY.

OpenStudy (zarkon):

in order for \[\int\limits \limits_{-\infty}^{\infty} x^{3}dx\] to converge the folowwing have to each converge \[\int\limits \limits_{-\infty}^{a} x^{3}dx+\int\limits \limits_{a}^{\infty} x^{3}dx\]

OpenStudy (anonymous):

BUT STILL IT IS INFINITY NOT ZERO

OpenStudy (zarkon):

it is not infinity...the integral does not converge

OpenStudy (zarkon):

alexray19, I believe, is under the false impression that \[\int\limits \limits_{-\infty}^{\infty} x^{3}dx\] is the same as \[\lim_{a\to\infty}\int\limits \limits_{-a}^{a} x^{3}dx\] which it is not.